OpenStudy (anonymous):

Write an appropriate direct variation equation if y = 18 when x = 2

OpenStudy (anonymous):

y varies directly as x. y = kx 5 = k * -10 k = -1/2 Therefore, y= -x/2 When x = 1, y = -1/2.